# Barrow entropy and spacetime foam

### v1

@@ I Introduction
As an intermediate stage of the transition from continuous to discrete space-time, we can consider the transformation of smooth surfaces into fractal ones. Even on the smallest scales currently observable, spacetime appears smooth and structureless. However, the opinion that continuous space-time is just an idealization is gradually winning out. One of the main reasons forcing us to abandon the concept of continuum is the existence of quantum fluctuations. It is natural to assume that quantum effects on Planck scales transform continuous space-time into a complex (turbulent) structure, called space-time foam (quantum foam) [6].
@@ IV Barrow entropy and spacetime foam
Barrow [16] proposed a phenomenological model of the formation of the fractal structure of the horizon for a stationary black hole. Burrow’s reasoning was as follows. The surface area of a Schwarzschild black hole $A_{g}=4\pi r_{g}^{2}$ determines its the entropy $S=A_{g}\frac{c^{3}}{4G}\approx\frac{A_{g}}{A_{Pl}}$. The entropy is proportional to the number of Planck squares on the surface of the horizon. Consequently, quantum gravitational effects leading to the fractal structure of the horizon can significantly increase its entropy compared to the standard Bekenstein entropy. Of course, the increase in the surface area of a black hole during its evolution is consistent with the second law of thermodynamics.
@@
where $A$ is the surface area of the horizon, $A_{P}$ is the Planck area, and $\Delta$ is a free parameter of the model $0\leq\Delta\leq 1$. In the undeformed case $(\Delta=0)$, the Burrow entropy coincides with the Bekenstein entropy, and the value $\Delta=1$ corresponding to the maximum deformation, leading to an increase in the fractal dimension of the horizon surface by one
@@
From here, we will find generalization of the fundamental inequality (2) for the Burrow entropy
@@
Burrow’s concept of entropy can be extended to cosmological horizons by introducing holographic dark energy with density $\rho\sim H^{2-\Delta}$ [17, 18, 19, 20, 21]. In this case, the parameter $\Delta$ can be expressed through the current values of cosmographic parameters $q_{0}j_{0}$ [22, 23]
@@
Note that in SCM $j=1$, therefore $\Delta=2$, that goes beyond the permissible values $\Delta$. In SCM, dark energy is realized in the form of a cosmological constant, and is naturally $\rho_{e}=CH^{2-\Delta}$ for $\Delta=2$ transformed into a constant. In this case entropy (it cannot be called Burrow entropy, for which $0\leq\Delta\leq 1$) $S_{B}=\left(\frac{A}{A_{0}}\right)^{1+\frac{\Delta}{2}}$ [24, 25]. Note that the Barrow entropy represents a special case of the 4-parameter construction that can reproduce the range of entropies used for horizon thermodynamics [26, 27].
@@ V Application examples
As Burrow noted [16], based on an analysis of the process of evaporation of black holes, the deformation of the horizon by quantum fluctuations should accelerate the evaporation process. We will consider this statement from a more general point of view.In its meaning, a time interval measured with a given accuracy represents the time of correct functioning of a clock or, in other words, its lifetime. Therefore, using inequality (3) we come to a far from obvious conclusion: the more accurate the clock, the less time it operates with a given accuracy. Let’s test this concept using a black hole with Bekenstein entropy ($\Delta=0$) as a clock. Then the accuracy $\delta t$ of such a device naturally choose $r_{g}/c$. Now, using the upper limit of inequality (3), we determine the lifetime of such a device (i.e., black hole)
@@
The resulting lifetime of a black hole in this unusual way coincides with the evaporation time of a black hole, which Hawking calculated. The only difference is in the numerical factor. Turning to Burrow entropy, we note that an increase in the degree of fractality (an increase in the parameter $\Delta$) leads to a decrease $\delta t$ ( increase in measurement accuracy). This, in turn, due to inequality (20), will reduce the time interval available for measurement, i.e. lifetime of the object (see Fig. 2).

### v2

@@ I Introduction
As an intermediate stage of the transition from continuous to discrete space-time, we can consider the transformation of smooth surfaces into fractal ones [6, 7]. Even on the smallest scales currently observable, spacetime appears smooth and structureless. However, the opinion that continuous space-time is just an idealization is gradually winning out. One of the main reasons forcing us to abandon the concept of continuum is the existence of quantum fluctuations. It is natural to assume that quantum effects on Planck scales transform continuous space-time into a complex (turbulent) structure, called space-time foam (quantum foam) [8].
@@ IV Barrow entropy and spacetime foam
Barrow [18] proposed a phenomenological model of the formation of the fractal structure of the horizon for a stationary black hole. Barrow’s reasoning was as follows. The surface area of a Schwarzschild black hole $A_{g}=4\pi r_{g}^{2}$ determines its the entropy $S=A_{g}\frac{c^{3}}{4G}\approx\frac{A_{g}}{A_{Pl}}$. The entropy is proportional to the number of Planck squares on the surface of the horizon. Consequently, quantum gravitational effects leading to the fractal structure of the horizon can significantly increase its entropy compared to the standard Bekenstein entropy. Of course, the increase in the surface area of a black hole during its evolution is consistent with the second law of thermodynamics.
@@
where $A$ is the surface area of the horizon, $A_{P}$ is the Planck area, and $\Delta$ is a free parameter of the model $0\leq\Delta\leq 1$. In the undeformed case $(\Delta=0)$, the Barrow entropy coincides with the Bekenstein entropy, and the value $\Delta=1$ corresponding to the maximum deformation, leading to an increase in the fractal dimension of the horizon surface by one
@@
From here, we will find generalization of the fundamental inequality (2) for the Barrow entropy
@@
Barrow’s concept of entropy can be extended to cosmological horizons by introducing holographic dark energy with density $\rho\sim H^{2-\Delta}$ [19, 20, 21, 22, 23, 24]. In this case, the parameter $\Delta$ can be expressed through the current values of cosmographic parameters $q_{0}j_{0}$ [25, 26, 27]
@@
Note that in SCM $j=1$, therefore $\Delta=2$, that goes beyond the permissible values $\Delta$. In SCM, dark energy is realized in the form of a cosmological constant, and is naturally $\rho_{e}=CH^{2-\Delta}$ for $\Delta=2$ transformed into a constant. In this case entropy (it cannot be called Barrow entropy, for which $0\leq\Delta\leq 1$) $S_{B}=\left(\frac{A}{A_{0}}\right)^{1+\frac{\Delta}{2}}$ [28, 29]. Note that the Barrow entropy represents a special case of the 4-parameter construction that can reproduce the range of entropies used for horizon thermodynamics [30, 31].
@@ V Application examples
As Barrow noted [18], based on an analysis of the process of evaporation of black holes, the deformation of the horizon by quantum fluctuations should accelerate the evaporation process. We will consider this statement from a more general point of view.In its meaning, a time interval measured with a given accuracy represents the time of correct functioning of a clock or, in other words, its lifetime. Therefore, using inequality (3) we come to a far from obvious conclusion: the more accurate the clock, the less time it operates with a given accuracy. Let’s test this concept using a black hole with Bekenstein entropy ($\Delta=0$) as a clock. Then the accuracy $\delta t$ of such a device naturally choose $r_{g}/c$. Now, using the upper limit of inequality (3), we determine the lifetime of such a device (i.e., black hole)
@@
The resulting lifetime of a black hole in this unusual way coincides with the evaporation time of a black hole, which Hawking calculated. The only difference is in the numerical factor. Turning to Barrow entropy, we note that an increase in the degree of fractality (an increase in the parameter $\Delta$) leads to a decrease $\delta t$ ( increase in measurement accuracy). This, in turn, due to inequality (20), will reduce the time interval available for measurement, i.e. lifetime of the object (see Fig. 2).